# Semantic-Aware Adaptive Video Streaming Using Latent Diffusion Models for Wireless Networks

### v1

@@ Abstract
This paper proposes a novel framework for real-time adaptive-bitrate video streaming by integrating latent diffusion models (LDMs) within the FFmpeg techniques. This solution addresses the challenges of high bandwidth usage, storage inefficiencies, and quality of experience (QoE) degradation associated with traditional constant bitrate streaming (CBS) and adaptive bitrate streaming (ABS). The proposed approach leverages LDMs to compress I-frames into a latent space, offering significant storage and semantic transmission savings without sacrificing high visual quality. While it keeps B-frames and P-frames as adjustment metadata to ensure efficient video reconstruction at the user side, the proposed framework is complemented with the most state-of-the-art denoising and video frame interpolation (VFI) techniques. These techniques mitigate semantic ambiguity and restore temporal coherence between frames, even in noisy wireless communication environments. Experimental results demonstrate the proposed method achieves high-quality video streaming with optimized bandwidth usage, outperforming state-of-the-art solutions in terms of QoE and resource efficiency. This work opens new possibilities for scalable real-time video streaming in 5G and future post-5G networks.
@@ I Introduction
In today’s digital age, the demand for streaming services with high-quality video content has surged alongside the proliferation of Internet-enabled devices, bridging the gap between media products and edge users [1]. Streaming services, which cater to a vast audience with diverse preferences, face the challenge of delivering superior video quality while optimizing wireless network bandwidth and resource usage. Real-time adaptive bit-rate streaming (ABS) has emerged as a vital technology in this landscape, which enables dynamically adjusting video quality based on network conditions to provide a seamless viewing experience and improved quality of experience (QoE) [2]. One significant advancement in ABS is the integration of fast forward moving picture experts group (FFmpeg), advanced video coding (H.264), hypertext transfer protocol (HTTP), and dynamic adaptive streaming over HTTP (DASH). These open-source multimedia frameworks are capable of decoding, encoding, transmission, and streaming digital media files in various formats. In particular, FFmpeg’s flexibility and efficiency make it an indispensable tool for real-time video processing and adaptive streaming solutions.
Diffusion models (DMs) [3], a class of Artificial Intelligence generative content (AIGC) models, have shown promising results in generating high quality media content. These models can also intelligently refine video streams at the receiver side, significantly improving visual quality while minimizing bandwidth requirements. By integrating latent diffusion models (LDMs) [4] into the FFmpeg framework, high-quality real-time video streaming can be achieved while dynamically adapting to varying network conditions and diverse quality of experience (QoE) requirements based on different media content. Along another note, the advent of fifth-generation (5G) and beyond 5G (B5G) networks offer unprecedented speeds and low latency, providing an ideal infrastructure for deploying sophisticated video streaming technologies. Despite their advantages, 5G and B5G networks may introduce more semantic ambiguities due to encoding techniques that are more sensitive to wireless channel noises, large-scale high-quality 4K/6K video streaming delivery, and the increased utilization of machine learning models, all of which must be addressed to the quality of the transmitted video. As demonstrated by the excellent performance of LDM in wireless image transmission and video generation tasks, advanced data compression and channel denoising techniques can also ensure a flawless streaming experience with higher QoE and quality of service (QoS) in wireless video streaming.
To this end, this paper first highlights the key challenges in the domain of conventional video streaming and DM-enabled video streaming. Next, we propose a novel LDM-enabled semantic-aware adaptive video streaming framework that enables transfer of high-quality video contents over wireless channels, considering different frame types  [^1] To efficiently transmit video within limited bandwidth, video streaming employs frame decomposition into three frame types: I-frames, P-frames, and B-frames [5]. The details are given in Section II-C. . Specifically, the proposed video streaming scheme integrates LDM and FFmpeg for efficient compression, reducing bandwidth by encoding keyframes as latent features and motion data as metadata. To enhance robustness, LDMs denoise I-frames and regenerate B/P-frames, ensuring better video quality over noisy wireless channels. Moreover, an adaptive bitrate mechanism using CNN-GRU optimizes streaming based on varying network conditions, media content, and user preferences for an improved viewing experience. The efficiency and reliability of the proposed system are quantified in terms of QoE that incorporates the trade-off between streaming video quality, quality fluctuation, and the risk of rebuffering events[6]. providing concrete evidence of its superiority over the other ABS algorithms such as, BOLA, Comyco, MERINA [6].
@@
The rest of this paper is organized as follows. Section II introduces existing challenges of video streaming and DM-aided semantic video communication. Section IV presents the design and workflow of the proposed LDM-enabled ABS framework, including its key components for compression, transmission, and video reconstruction. Section V provides numerical results and performance analysis to validate the effectiveness of the proposed approach in real-time streaming scenarios. Finally, Section VI concludes the paper and outlines future research directions.
II Fundamental Challenges in Semantic Communication for Video Streaming
This section outlines several fundamental challenges in semantic-aware video streaming, including constant bit-rate streaming (CBS), adaptive bit-rate streaming (ABS), and DM-enabled bit-rate streaming. A comparative analysis of the various bit-rate streaming scheme is presented in Fig. 1. In CBS, parallel channels are utilized for transmission within the wireless zone to the user side, requiring extensive wireless resources. In ABS, a single channel is transmitted to the user side, and frames can be downscaled in the computing zone based on the user’s request. This approach conserves wireless transmission resources while providing adaptive bitrate options. LD-ABS introduces a novel mechanism where only latent frames are transmitted, eliminating the need for complete frames. This design significantly compresses the transmission, reducing bandwidth and storage requirements, and enabling ultra-efficient streaming.
@@ II-A High Bandwidth and Storage Utilization for CBS
Multi-bitrate streaming consumes a substantial amount of bandwidth and resources for transmission. To satisfy the diverse customers’ demands of streaming video, traditionally providers adopt constant bitrate streaming (CBS), which involves providing parallel streaming channels to broadcast video in different resolutions and qualities.
(i) Bandwidth Consumption: CBS requires a fixed amount of bandwidth for each resolution and quality level, irrespective of the content’s complexity [7]. This can lead to inefficient use of available bandwidth, especially during periods of low network traffic or when streaming less complex scenes [8]. CBS necessitates maintaining multiple streams for different quality levels, which can overload the content delivery network (CDN) [6]. In the context of CBS, as illustrated in Fig. 1, let’s assume we have 4 parallel routes to stream videos. We capture video at a 1080P resolution source. Each route streams video at 1080P, 720P, 360P, and 144P resolutions, respectively. FFMpeg decodes the video with different bitrates to their respective reference frames $I_{n}$ and affiliated predictive frames $P_{n}$ and $B_{n}$. Note that the size of predictive frames is significantly smaller than the reference frames [9].
Users manually select the resolution on the player side. This design requires all frames across all resolutions to be sent to the user side. Broadcasting video in different bitrates consumes considerable traffic and resources, overburdening the CDN. Overall, CBS utilizes the highest storage and resources due to multi-bitrate streaming.
(ii) Server Overload: Handling numerous simultaneous requests for high-quality video streams strain server resources and place a heavy burden on networking infrastructure and cloud resources, leading to increased latency [7], buffering issues [8], and server crashes at peak times or major events.
(iii) Wireless Channel Impairments: In wireless communication systems, the variability in signal quality and network conditions makes it difficult to maintain a constant bitrate. High-bandwidth streams can quickly exhaust available wireless resources, leading to degraded QoE for mobile users and increased traffic usage [7].
@@
II-B Low QoE Multi-Bitrate Adaptive Streaming
@@
ABS dynamically adjusts video quality based on a user’s network conditions and device capabilities to optimize viewing experiences while efficiently utilizing bandwidth. Unlike CBS, which delivers pre-defined quality streams through parallel channels, ABS continuously monitors real-time factors such as network throughput, buffer occupancy, and device performance to make fine-grained adjustments. Achieving high-quality video streaming without excessive bandwidth consumption remains a challenging task. Traditional ABS algorithms often struggle to balance video quality and network efficiency, leading to either compromised video quality or excessive data usage [2]. Video streaming services require ABS to cater to varying network conditions and device capabilities.
(i) Challenges of diverse wireless conditions: Compared to CBS, which does not automatically select chunks, ABS must dynamically adjust the video bitrate in response to varying network conditions. This adaptation introduces potential latency during bitrate switching, especially in dynamic wireless environments. Additionally, ABS aggressively reduces bitrate under poor network conditions, often resulting in compression artifacts. In contrast, CBS maintains a consistent quality, reducing the risk of rebuffering but lacking the flexibility to optimize for fluctuating bandwidth. [6].
(ii) Cache and Buffering Issues: Managing cache and buffering of short clips in ABS streaming is complex. Storing multiple resolution video clips adds complexity and consumes significant storage and bandwidth[7].
(iii) Downscaling to Ensure Users Watching Smooth and Continuous Video: Existing ABS methods attempt to maintain smooth play and playback by frequently switching bitrates. However, excessive bitrate switching negatively impacts QoE and QoS for viewers [7].
II-C VFI and Temporal Consistency
@@
Streaming video is typically decomposed into Intra-coded frames (I-frames), Predictive-coded frames (P-frames) and Bidirectionally predictive-coded frames (B-frames) [5], which are defined, respectively, as follows:
  • I-frames are compressed using only the information within the frame itself. They contain a complete image, similar to a PNG image file.
@@
  • P-frames use data from the previous I-frame or P-frame to predict and encode only the differences (motion vectors and residual data).
  • B-frames use data from both preceding and following I-frames or P-frames to predict the frame. They encode differences from both directions, often achieving higher compression.
Reconstructing frames based on these reference frames and adjustment frames is difficult [10]. In video compression, I-frames serve as key reference frames containing the full image data, while P- and B-frames store only the motion vectors [^2] Motion vectors capture the displacement of objects between frames, encoding movement without redundantly storing unchanged visual information already present in the I-frame. and differences relative to I-frames. Motion vectors are light compared to frames, allowing for efficient reconstruction while maintaining high compression efficiency. However, estimating motion vectors between frames is complex. Ensuring temporal consistency across frames is challenging, especially when using generative models for frame reconstruction [10]. In addition, accurate motion estimation is crucial for aligning the reconstructed P-frames and B-frames with the original video sequence. Errors in motion estimation can result in misalignment and temporal inconsistencies [10].
II-D High Latency and Low Resolution Restoration on Reverse Diffusion Frame Reconstruction
Traditional diffusion models require multiple steps to generate high-resolution frames, making them impractical for real-time video applications, which are restricted to 500ms [4].
(i) Slow reconstruction speed: The reverse diffusion process involves gradually denoising the corrupted frames by iteratively applying the diffusion model which result in high latency, especially for long video sequences [11, 10, 12].
(ii) Computational complexity: Reconstructing all the frames across multiple resolutions overburdens VAE.
(iii) Limited frame generalization: DMs are trained on specific datasets and may not generalize well to unseen or diverse content. In addition, streaming video contains semantic errors. These difficulties lead to challenges on restoring low-resolution frames from different sources or with different characteristics [11].
III State-of-the-Art: DM-Aided Video Streaming
Traditional video compression standards, such as H.264, present challenges in terms of bandwidth consumption, storage requirements, and computational overhead [9]. To achieve real-time performance, dedicated hardware acceleration (e.g., specialized GPUs or ASICs) is often required for multi-channel H.264 encoding to prevent frame drops and buffering delays on the producer side [13]. Additionally, packet loss in predicted frames (P- or B-frames) can degrade following frames until an I-frame refreshes the stream. To address the aforementioned issues, LDMs and DMs are becoming popular. In this section, we introduce the fundamental principles of LDMs and compare the key differences between DMs and LDMs followed by a review of existing literature in the domain of DM/LDM-enabled video streaming.
III-A Overview of Latent Diffusion Models
LDMs specifically operate in a lower-dimensional latent space rather than pixel space, leading to more efficient computation and better scalability [4]. Conditional LDMs (CLDMs) further enhance this approach by conditioning the diffusion process on additional information, such as motion adjustment metadata. Among various implementations, stable diffusion [3] achieve high-quality outputs with improved efficiency and reduced computational requirements. Stable diffusion’s architecture leverages both LDMs and CLDMs, allowing it to generate detailed and coherent frames while maintaining a smaller memory footprint compared to other models. This makes LDMs particularly suitable for compressing streaming video, as they can effectively balance the trade-offs between computational load, memory usage, and output quality, enabling efficient and scalable video compression. The key distinctions of LDMs compared to traditional DMs are:
$\bullet$ Restoration of original data: LDMs learn to restore the original data by reversing the noising process through incremental denoising and reconstruction.
$\bullet$ Systematic degradation of training data: Both LDMs and DMs systematically introduce Gaussian noise to degrade the original data in a step-by-step process known as diffusion. This transformation simplifies the data distribution.
$\bullet$ Operating in latent space: Unlike traditional DMs that operate directly in pixel space, LDMs function in a lower-dimensional semantic latent space. This shift leads to more efficient computation and improved scalability.
$\bullet$ Conditional diffusion: CLDMs enhance the standard diffusion process by incorporating additional information, such as motion adjustment metadata, to condition the model denoising processes. This results in higher quality and more relevant generated frames.
$\bullet$ Denoising via Neural Network: LDMs utilize neural networks to capture complex spatiotemporal relationships across different frames in video streaming, facilitating the generation of high-fidelity frames and enhancing video synthesis.
III-B Recent Advancements on DM-aided Video Streaming
III-B1 DM-aided Video Streaming

### v2

@@ Abstract
This paper proposes a novel Semantic Communication (SemCom) framework for real-time adaptive-bitrate video streaming by integrating Latent Diffusion Models (LDMs) within the FFmpeg techniques. This solution addresses the challenges of high bandwidth usage, storage inefficiencies, and quality of experience (QoE) degradation associated with traditional Constant Bitrate Streaming (CBS) and Adaptive Bitrate Streaming (ABS). The proposed approach leverages LDMs to compress I-frames into a latent space, offering significant storage and semantic transmission savings without sacrificing high visual quality. While retaining B-frames and P-frames as adjustment metadata to support efficient refinement of video reconstruction at the user side, the proposed framework further incorporates state-of-the-art denoising and Video Frame Interpolation (VFI) techniques. These techniques mitigate semantic ambiguity and restore temporal coherence between frames, even in noisy wireless communication environments. Experimental results demonstrate the proposed method achieves high-quality video streaming with optimized bandwidth usage, outperforming state-of-the-art solutions in terms of QoE and resource efficiency. This work opens new possibilities for scalable real-time video streaming in 5G and future post-5G networks.
@@ I Introduction
In today’s digital age, the demand for streaming services with high-quality video content has surged alongside the proliferation of Internet-enabled devices, bridging the gap between media products and edge users [1]. Streaming services, which cater to a vast audience with diverse preferences, face the challenge of delivering superior video quality while optimizing wireless network bandwidth and resource usage. Real-time Adaptive Bit-rate Streaming (ABS) has emerged as a vital technology in this landscape, which enables dynamically adjusting video quality based on network conditions to provide a seamless viewing experience and improved Quality of Experience (QoE) [2]. One significant advancement in ABS is the integration of Fast Forward Moving Picture Experts Group (FFmpeg), advanced video coding (H.264), Hypertext Transfer Protocol (HTTP), and Dynamic Adaptive Streaming over HTTP (DASH). These open-source multimedia frameworks are capable of decoding, encoding, transmission, and streaming digital media files in various formats. In particular, FFmpeg’s flexibility and efficiency make it an indispensable tool for real-time video processing and ABS solutions.
Compared with traditional video-streaming technologies, Semantic Communication (SemCom) provides a next-generation paradigm for more efficient, robust, and semantic-aware video transmission that addresses many limitations of conventional video streaming. Diffusion Models (DMs) [3], a class of Artificial-Intelligence Generative-Content (AIGC) models, have demonstrated a remarkable ability to generate high-quality media. In a SemCom system, such models can be used for efficient encoding, robust denoising, and high-fidelity decoding, enabling intelligent fine-tuning of incoming video streams at the receiver and thereby improving both perceptual and semantic integrity of the reconstructed video frames. Fifth-generation (5G) and Beyond-5G (B5G) networks deliver unprecedented throughput and ultra-low latency, forming an ideal foundation for DM-enabled video streaming. However, these networks impose greater resource demands, introduce semantic ambiguities through encoding and decoding schemes that are highly sensitive to wireless channel noise, large-scale 4K and 6K streaming requirements, and the extensive use of machine-learning models, all of which must be managed to preserve end-to-end video quality. Fortunately, Latent Diffusion Models (LDMs) [4] have achieved excellent performance in wireless image transmission and video generation, and their advanced compression and channel-denoising techniques can outperform pixel-space DMs, ensuring seamless streaming with higher QoE and Quality of Service (QoS).
Inspired by both conventional FFmpeg video-streaming methods and emerging LDM-enabled SemCom for video transmission, we integrate LDM [4] into the FFmpeg framework to achieve high-quality real-time streaming that adapts dynamically to network fluctuations and diverse QoE requirements. Specifically, this paper first surveys the key challenges in conventional video transmission, SemCom-based approaches, and DM-enabled streaming. Next, we introduce an LDM-enabled semantic-aware ABS framework capable of delivering high-quality content over wireless channels by leveraging frame-type information [^1] To efficiently transmit video within limited bandwidth, streaming standards decompose video into I-frames, P-frames, and B-frames [5].. The proposed scheme compresses I-frames into low-dimensional latent representations and encodes P and B frame motion as compact metadata, achieving significant bandwidth savings. At the receiver, LDMs denoise the semantic latent space of I-frames and refine the latent representations of P-frames and B-frames to enhance robustness against wireless channel noises. A Convolutional Neural Network–Gated Recurrent Unit (CNN–GRU)-based adaptive-bitrate controller further optimizes streaming decisions according to real-time network conditions, content characteristics, and user preferences. We evaluate the system using a comprehensive QoE metric that balances video quality, quality fluctuations, and rebuffering risk, and demonstrate its superiority over state of the art ABR algorithms, including BOLA, Comyco, and MERINA [6].
@@
The rest of this paper is organized as follows. Section II introduces the existing challenges and limitations of conventional video streaming and SemCom for video transmission. Section III provides a brief overview of DMs and LDMs, discusses SemCom-based video transmission methods built upon these models, and highlights the remaining open issues. Section IV presents the design and workflow of the proposed LDM-enabled ABS framework, detailing its key components for compression, transmission, and video reconstruction. Section V provides numerical results and performance analysis to validate the effectiveness of the proposed approach under real-time streaming scenarios. Finally, Section VI concludes the paper and outlines potential future research directions.
II Fundamental Challenges in SemCom for Video Streaming
This section first outlines three major challenges in traditional video streaming: high bandwidth and storage consumption (Subsection II-A), low QoE (Subsection II-B), and Video Frame Interpolation (VFI) (Subsection II-C). It then introduces the concept of SemCom and how it addresses these challenges in Subsection II-D.
@@ II-A High Bandwidth and Storage Utilization for CBS
Multi-bitrate streaming consumes substantial bandwidth and transmission resources. To meet diverse QoE demands, providers traditionally adopt Constant Bitrate Streaming (CBS), as depicted in Fig. 1, wherein parallel streaming channels deliver videos at multiple resolutions and quality levels. The primary challenges associated with CBS are outlined as:
(i) Bandwidth Consumption: CBS requires allocating a fixed bandwidth for each resolution and quality level, regardless of the video’s content complexity [7]. This results in inefficient utilization of available bandwidth, especially during periods of low network congestion or when transmitting less complex scenes [8]. As shown in Fig. 1, a typical CBS system maintains four parallel streams at 1080P, 720P, 360P, and 144P resolutions, respectively. Videos captured at 1080P are transcoded into different bitrates using FFmpeg, generating corresponding reference frames ($I_{n}$) and predictive frames ($P_{n}$ and $B_{n}$), where predictive frames are significantly smaller in size [9]. However, all frames across all resolutions must still be transmitted to users, as they manually select the desired resolution at the client side. This redundancy leads to substantial traffic and storage overhead, overburdening the Content Delivery Network (CDN) and causing excessive resource consumption.
(ii) Server and Infrastructure Overload: CBS places a heavy burden on servers, networking infrastructures, and cloud platforms. Managing numerous simultaneous requests for high-quality streams increases processing pressure, exacerbates network congestion, and leads to higher latency, frequent buffering, and even service disruptions during peak periods or large-scale events [7, 8]. The need to sustain multiple concurrent streams for various resolutions further amplifies server-side storage and computation demands.
(iii) Wireless Transmission Challenges: In wireless communication environments, maintaining constant bitrates becomes highly challenging due to the inherent variability of signal quality and network conditions. High-bitrate streams under CBS rapidly deplete available wireless resources, leading to degraded QoE for mobile users, increased retransmissions, and higher traffic costs [7].
@@
II-B Low QoE in Multi-Bitrate Adaptive Streaming
@@
ABS dynamically adjusts video quality according to users’ network conditions and device capabilities, aiming to optimize viewing experiences while efficiently utilizing bandwidth. However, achieving high-quality video streaming without excessive bandwidth consumption remains challenging. Traditional ABS approaches often struggle to balance video quality and network efficiency, resulting in either compromised viewing experiences or unnecessary data consumption [2]. The key challenges are summarized as follows:
(i) Adaptation under Variable Wireless Conditions: ABS must dynamically adjust the video bitrate in response to fluctuating network environments. Frequent bitrate switching can introduce latency and instability, particularly in dynamic wireless scenarios. Under poor conditions, ABS tends to aggressively lower the bitrate, causing compression artifacts and degraded visual quality [6], whereas CBS maintains a consistent quality level, albeit with limited flexibility.
(ii) Cache and Buffer Management: Managing cached video chunks at multiple resolutions increases system complexity and storage demands [7]. Efficiently prefetching, storing, and switching between segments remain a critical challenge for delivering seamless playback in ABS systems.
(iii) Trade-off Between Smooth Playback and Quality Consistency: To ensure smooth and continuous playback, ABS frequently switches bitrates according to network variations. However, excessive bitrate fluctuation negatively impacts both QoE and QoS [7], leading to inconsistent viewing quality and user dissatisfaction.
II-C Video Frame Interpolation and Temporal Consistency
@@
Streaming video is typically composed of Intra-coded frames (I-frames), Predictive-coded frames (P-frames), and Bidirectionally predictive-coded frames (B-frames) [5], defined as follows:
  • I-frames: Independently compressed frames containing complete image information, similar to standalone images (e.g., PNG files).
@@
  • P-frames: Frames that predictively encode differences from preceding I-frames or P-frames using motion vectors and residual data.
  • B-frames: Frames that predictively encode differences by referencing both preceding and following I-frames or P-frames, offering higher compression efficiency.
Reconstructing video sequences from these frames is inherently challenging [10]. In video compression, I-frames serve as key references containing full image data, while P-frames and B-frames encode only motion vectors [^2] Motion vectors represent object displacements between frames, enabling efficient compression by avoiding redundant storage of unchanged content. and residual differences relative to reference frames. Although motion vectors are lightweight compared to full frames, enabling high compression ratios, accurate motion estimation between frames is difficult. Temporal consistency is particularly hard to maintain when using generative models for frame reconstruction [10]. Moreover, errors in motion estimation can cause misalignment and temporal artifacts, leading to QoE degradation [10].
II-D SemCom for Video Streaming and Challenges
SemCom focuses on transmitting the meaning of source data rather than raw bits or symbols. By extracting, encoding, and adapting semantic features, SemCom can directly address some challenges in traditional video transmissions, i.e.,
(i) SemCom reduces bandwidth and storage: SemCom encoders extract only the essential semantic features of each video segment, eliminating parallel full-resolution streams and cutting redundant data. This feature-level compression sharply lowers both bandwidth use and storage requirements.
(ii) Semantic-guided caching, buffering, and reconstruction: Edge or client caches store compact semantic descriptors instead of full video segments, simplifying buffer management and speeding startup. At playback, generative decoders use these semantic cues to rebuild frames with strong temporal coherence and reduced interpolation artifacts.
Nevertheless, SemCom faces several challenges that motivate the use of AIGC models: (i) Limited semantic reconstruction fidelity: SemCom systems rely on light-weight models that struggle to reproduce fine details, complex textures, and rapid motions, resulting in blurring, artifacts, and temporal inconsistencies under demanding scenarios. (ii) Wireless impairments and noise mitigation: Most SemCom system designs assume ideal or simplified channels and lack integrated denoising mechanisms, leaving semantic features vulnerable to real-world impairments such as multipath fading and interference.
III State-of-the-Art: DM-Aided SemCom for Video Streaming
To address the key challenges currently focused on in SemCom for video streaming, DMs and LDMs, as the most advanced AIGC models for visual data, offer effective solutions for high-quality data reconstruction and robust information transmission in SemCom. Therefore, this section first briefly reviews DMs and LDMs in Subsection III-A, then introduces several DM- and LDM-based approaches applied to video streaming in Subsection III-B, and finally summarizes the remaining limitations of these methods in Subsection III-C.
III-A Overview of DMs and LDMs
DMs are a class of likelihood-based generative frameworks that learn to synthesize data by reversing a controlled noising process. During training, each data sample is gradually corrupted through a sequence of Gaussian noise steps; a neural network then learns to denoise these noisy samples step by step, ultimately recovering the original signal. While this pixel-space diffusion yields high-quality outputs, it demands substantial computation and memory, especially for high-resolution video. Compared to DMs, LDMs address these challenges by shifting both the forward noise injection and reverse denoising into a lower-dimensional latent space. An encoder first projects each video frame into compact latent embeddings; diffusion and denoising occur entirely within this latent domain; a decoder then reconstructs the full-resolution frames [4]. Conditional LDMs (CLDMs) further enhance this process by integrating auxiliary information—such as motion vectors or semantic descriptors—into the denoising network. The key distinctions of LDMs compared to DMs for video streaming are:
  • Latent-space diffusion and dimensionality reduction: By encoding frames into a compact latent representation and performing diffusion steps in this lower-dimensional space, LDMs drastically reduce both memory usage and computational load, making high-resolution video generation more tractable.
  • Multi-step denoising with temporal coherence: The sequential denoising process refines latent embeddings incrementally, enabling fine-grained reconstruction. This multi-step approach captures complex spatial details and temporal dependencies across adjacent frames, allowing robust recovery from noisy or incomplete latents.
  • Conditioned Denoising: By conditioning the denoising network on external cues (such as motion adjustment metadata, scene semantics, or object trajectories), CLDMs guide the generation toward temporally consistent and semantically accurate video frames even under varying network or content dynamics.
III-B Existing State-of-the-Art
III-B1 DM-aided SemCom for Video Streaming